Due to erratic sales of its sole product-a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Req3 Reg 4 Req 5A Req 5B Req 5C Compute the coReg 1 Reg 2 Req3 Reg 4 Req 5A Reg 5B Req 5C The president believes that a $6,800 increase in the monthly advertising budget,Reg 1 Req 2 Req3 Reg 4 Req 5A Reg 5B Reg 5C Refer to the original data. The sales manager is convinced that a 10% reduction iReg 1 Reg 2 Req3 Reg 4 Reg 5A Reg 5B Req 5C Refer to the original data. The Marketing Department thinks that a fancy new packReq 1 Reg 2 Reg 3 Reg 4 Reg 5A Req 5B Req 5C Refer to the original data. By automating, the company could reduce variable expReg 1 Reg 2 Reg 3 Reg 4 Req 5A Req 5B Req 5C Refer to the original data. By automating, the company could reduce variable expReg 1 Reg 2 Reg 3 Req 4 Req 5A Req 5B Reg 5C Refer to the original data. By automating, the company could reduce variable exp
Answer #1
Req 1 :
Contribution margin per unit = Contribution margin / Units sold = 198000 / 13200 = 15
CM ratio = Contribution margin / Sales = 198000 / 396000 = 50%
Break-even point in units = Fixed expenses / Contribution margin per unit = 220500 / 15 = 14700
Break-even point in dollars = Fixed expenses / Contribution margin ratio = 220500 / 50% = 441000
Req 2 :
Increase in Contribution margin ( Increase in sales * Contribution margin ratio = 88000 * 50% ) 44000
(-) Increase in advertising budget 6800
Increase (decrease) in net operating income 37200
Net operating income increases by 37200
Req 3 :
Revised selling price = Current selling price * ( 1 - % reduction ) = 30 * ( 1 - 10% ) = 27
Current unit variable cost = Variable expenses / Units sold = 198000 / 13200 = 15
Revised fixed costs = Current fixed costs + Increase in advertising expense = 220500 + 31000 = 251500
Revised sales units = Current sales units * 2 = 13200 * 2 = 26400
Sales ( 26400*27 ) 712800
(-) Variable expenses ( 26400*15 ) 396000
Contribution margin 316800
(-) Fixed expenses 251500
Net operating Income (loss) 65300
Revised net operating Income (loss) 65300
Req 4 :
Revised unit variable cost = Current unit variable cost + 0.70 = 15 + 0.70 = 15.70
Units sales to attain target profit = ( Target profit + Fixed expenses ) / ( Selling price - Unit variable cost ) = ( 4100 + 220500 ) / ( 30 - 15.70 ) = 15706
Req 5A :
Current unit variable cost = Variable expenses / Units sold = 198000 / 13200 = 15
Revised unit variable cost = 15 - 3 = 12
Revised fixed expenses = 220500 + 50000 = 270500
Contribution margin per unit = Selling price - Unit variable cost = 30 - 12 = 18
CM ratio = Contribution margin per unit / Selling price = 18 / 30 = 60%
Break-even point in unit sales = Fixed costs / Contribution margin per unit = 270500 / 18 = 15028
Break even point in dollar sales = Fixed costs / CM ratio = 270500 / 60% = 450833
Req 5B :
Not automated Automated
Total Per unit % Total Per unit %
Sales 624000 30.00 100.00% 624000 30.00 100.00%
Variable expenses 312000 15.00 50.00% 249600 12.00 40.00%
Contribution margin 312000 15.00 50.00% 374400 18.00 60.00%
Fixed expenses 220500 270500
Net operating income 91500 103900
Req 5C :
Answer : Yes
Explanation : If the company sells 20800 units, then the company will earn more net operating income if it automates its operations.